# OpenEarthAgent[Uncaptioned image]: A Unified Framework for Tool-Augmented Geospatial Agents

### v2

@@ Abstract
Recent progress in multimodal reasoning has enabled agents that interpret imagery, connect it with language, and execute structured analytical tasks. Extending these capabilities to remote sensing remains challenging, as models must reason over spatial scale, geographic structures, and multispectral indices while maintaining coherent multi-step logic. To address this gap, we introduce OpenEarthAgent, a unified framework for tool-augmented geospatial reasoning trained on satellite imagery, natural-language queries, and structured reasoning traces. Beyond serving as a benchmark, OpenEarthAgent establishes a cohesive agentic architecture built around a unified executable tool registry and trajectory-based policy learning. The framework standardizes heterogeneous visual, spectral, GIS, and georeferenced raster operations under a consistent callable schema, enabling modular orchestration and deterministic execution. Training is performed via supervised fine-tuning on structured reasoning trajectories, coupled with deterministic replay validation to ensure full-chain executability and spatial correctness. The accompanying corpus comprises 14,538 training and 1,169 evaluation instances with over 107K reasoning steps, spanning urban, environmental, disaster, and infrastructure domains and incorporating GIS-based operations alongside index analyses such as NDVI, NBR, and NDBI. Grounded in explicit multi-step traces, the learned agent exhibits structured reasoning, stable spatial grounding, and interpretable tool-driven behavior across diverse EO scenarios. We report consistent improvements over a strong baseline and competitive performance relative to recent open and closed-source models. Grounded in explicit reasoning traces, the learned agent demonstrates structured reasoning, stable spatial understanding, and interpretable behaviour through tool-driven geospatial interactions across diverse conditions. We report consistent improvements over a strong baseline and competitive performance relative to recent open and closed-source models.
@@ 2 Related Work
An emerging wave of EO research is now bringing these agentic principles into geospatial contexts. ThinkGeo [shabbir2025thinkgeo] frames EO question answering as tool-augmented reasoning, benchmarking ReAct-style chains and exposing persistent weaknesses in coordinate consistency, spatial grounding, and multi-step planning. Earth-Agent [feng2025earth] broadens the tool ecosystem to include spectral products and standardized interfaces but still relies on largely predefined workflows, limiting physically verifiable GIS and index-based reasoning. Geo-OLM [stamoulis2025geo] explores plan-tool-verify prompting for compact models, improving efficiency but depending primarily on prompt-level heuristics rather than learned adaptive policies. Building on this gap, we introduce an EO-native agentic stack that unifies dataset construction, supervised reasoning alignment, and evaluation under a physically grounded framework. OpenEarthAgent enables broad orchestration of GIS, index-based, optical, and SAR operations through structured instruction-tool pairs; trains on detailed reasoning traces with a dedicated evaluation set; and supports interpretable, verifiable, multi-step reasoning across diverse geospatial conditions.
@@ 3.1 Dataset Curation
Table 2: Comprehensive overview of data sources used for dataset construction, covering optical, SAR, GIS, and multispectral domains. These sources collectively offer complementary spatial resolutions, and annotation formats, ranging from urban and transport monitoring to environmental and disaster assessment, forming a rich foundation for multi-sensor, tool-based geospatial reasoning. GSD = ground sample distance, B-Box = bounding box, Seg. = segmentation.
@@
Figure 2: Unified data-curation pipeline for OpenEarthAgent Dataset. The process integrates diverse data sources: RGB-SAR imagery, GIS-based spatial layers, and index-driven datasets to build a harmonized geospatial corpus. Each branch performs its own filtering: RGB-SAR samples undergo sufficiency and annotation checks; GIS scenes are selected using point-of-interest (POI) constraints; and index-based inputs capture spectral changes (e.g., NDVI, NBR, NDBI). The outputs are merged into unified JSON records containing imagery metadata and geopackage information. A question synthesizer and conversational one-shot generator create tool-grounded queries and reasoning traces, which are validated through an automated suite ensuring geographic and syntactic correctness. Invalid instances are discarded, while verified ones form the final high-quality corpus used for agentic training and evaluation.
Figure 3: Representative examples of query-reasoning trajectories from the proposed dataset. Each example illustrates the interleaving of natural-language thoughts, tool executions, and intermediate observations across diverse modalities and task types, including object localization, distance measurement, spectral index mapping, change detection, and GIS-based spatial analysis. These samples depict simplified excerpts, while the full dataset contains longer, more comprehensive multi-step reasoning traces.
@@
Dataset Statistics: Fig. 4 summarizes the scale, distribution, and structure of the OpenEarthAgent dataset. The training split contains 14,538 instances, and the benchmark split contains 1,169 evaluation samples. Each instance includes multimodal imagery, a natural-language query, and a structured reasoning trace with interleaved tool calls and intermediate observations. Across the training set, the corpus contains 100,656 reasoning steps (average 6.92 per query)and the benchmark split contains 7,064 reasoning steps (average 6.04 per query). Reasoning steps correspond to explicit thought-action-observation transitions, while tool calls capture the diversity of operational pathways required for solving curated agentic tasks. The breadth of modalities, thematic coverage, and depth of reasoning traces enable rigorous evaluation of grounded, interpretable, and tool-augmented geospatial reasoning.
@@ 3.2 Methodology
3.2.1 Overview:
We present OpenEarthAgent, a unified agentic learning framework designed for multimodal reasoning over earth observation data. The framework integrates tool-driven interaction and language-grounded reasoning into a single end-to-end formulation. Unlike conventional multimodal models that process all inputs in one step, OpenEarthAgent decomposes geospatial reasoning into an organized sequence of perception and action steps, where the agent learns to perceive, reason, and act through callable tools operating on visual, textual, and spatial modalities. Given a natural-language instruction $\mathbf{u}$ and corresponding multimodal inputs $\mathbf{v}$ (e.g., RGB imagery, paired change-detection inputs, SAR, GIS layers, or georeferenced raster data), the agent generates multi-step reasoning trajectories that interleave internal thoughts with explicit tool invocations and their returned outputs (see Fig. 5). At each step, the agent maintains a short-term working memory that contains: instruction, observations, prior tool calls, spatial metadata, and contextual feedback from previous executions. This working memory enables iterative reasoning and consistent spatial grounding. Training follows a supervised fine-tuning procedure on validated reasoning traces, improving syntactic validity, spatial consistency, and multi-step reasoning coherence across diverse geospatial contexts.
3.2.2 Unified Tool Registry:
To standardize the diverse operators used in EO reasoning, we construct a unified tool registry that abstracts heterogeneous visual, spectral, and GIS operations through a consistent callable schema:
@@ 3.2.2 Unified Tool Registry:
To ensure efficiency and consistency, intermediate outputs (e.g., derived vector layers, raster subsets, index maps, computed geometries) are stored in a tool execution cache memory. This cache includes geospatial archives and vector/raster layers that can be reused by subsequent tool calls within the same trajectory. Caching prevents redundant computation and ensures deterministic replay of reasoning steps. New tools can be integrated by registering their schema $\mathcal{M}_{j}$ without retraining the backbone model, ensuring extensibility while maintaining interface consistency. We outline the tool categories below.
a) Perceptual Tools. These operators form the perceptual layer of the agent, grounding language into image-space entities. Functions such as TextToBbox, ObjectDetection, and RegionAttributeDescription convert free-form queries into spatial primitives such as bounding boxes and semantic tags. CountGivenObject and SegmentObjectPixels extend these capabilities toward quantitative perception and support object counting and pixel-level area estimation. ChangeDetection bridges spatial and temporal understanding by analyzing multi-date pairs to detect semantic or structural changes such as damage or flooding.
b) GIS Computation Tools. To enable explicit geographic reasoning, the registry includes tools for geometric and coordinate-aware computation. Operators such as GetAreaBoundary, AddPoisLayer, and ComputeDistance manipulate vector geometries within standard coordinate reference systems. Outputs include numerical measures (e.g., distances, buffers, zonal statistics) or derived layers that can be visualized via DisplayOnMap. These tools support verifiable geodesic reasoning and metric-scale analysis.
c) Spectral Tools. A subset of operators performs band arithmetic and spectral analysis using raster inputs. AddIndexLayer, ComputeIndexChange, and ShowIndexLayer compute and visualize vegetation, burn, or built-up indices (e.g., NDVI, NBR, NDBI). These tools expose physically meaningful variables and enable quantitative reasoning over environmental dynamics.
d) Georeferenced Raster Tools. Tools such as GetBboxFromGeotiff and DisplayOnGeotiff operate directly on georeferenced raster data. They allow coordinate extraction, bounding box derivation, and visualization aligned with spatial metadata. These operators ensure spatial consistency when reasoning over projected raster datasets.
e) Utility Tools. Generic reasoning utilities such as Calculator, Solver, and Plot support arithmetic computation and symbolic reasoning. Visualization aids like DrawBox and AddText render intermediate reasoning results. OCR and GoogleSearch extend multimodal understanding through text extraction and contextual retrieval. Terminate signals completion of a reasoning trajectory.
3.2.3 Reasoning Trajectories and Training Objective
Each training sample encodes a structured reasoning trajectory that connects an instruction $\mathbf{u}$, observations $\mathbf{v}$, and a sequence of tool interactions. Formally, the trajectory is represented as:
@@ 4 Experiments
In this section, we outline implementation details, describe the step-by-step and end-to-end evaluation modes, and define metrics used to quantify reasoning fidelity, tool effectiveness, and geospatial accuracy.
4.1 Implementation Details
OpenEarthAgent is finetuned using Qwen3-4B-Instruct-2507 as the base model. Training was performed on LLM controller on four NVIDIA A100 GPUs (40 GB). The base model is loaded via Unsloth’s FastLanguageModel, enabling efficient multi-GPU distributed training. The model is trained for one epoch with a learning rate of $2\times 10^{-5}$, cosine learning rate scheduling, 0.05 warmup ratio, and a batch size of 16. The maximum sequence length is set to 4096 tokens to support long-context conversational tool interactions.
LLM is finetuned on conventional data, where the model alternates between text generation and tool invocation, providing tool outputs and feedback in JSON format, embedded in its user and assistant turns. Training applies response-only masking, computing loss exclusively on assistant-generated tokens, enabling accurate tool calls and structured responses without optimizing on prompt text or external tool outputs. This enables the model to establish a good grounding for tool invocation.
@@ 4.2 Baseline Evaluation
End-to-End Evaluation: This mode tests full autonomous execution with live tool use. The model issues tool calls, forms arguments, and reasons step-by-step based on prior outputs. This setting captures operational robustness, error handling, and how well the model links perception with action in geospatial workflows.
@@
4.3 Results
In step-by-step evaluation (Table 3), frontier models lead, with o4-mini scoring highest on Summ. (89.48) and gpt4o showing strong Inst./Tool. accuracy. Among open-source models, Qwen2.5-Instruct (7B) performs competitively (Inst. 94.08, Tool. 85.51), whereas other 7B–8B models show notable degradation, particularly on argument value prediction. OpenEarthAgent (4B) substantially narrows this gap despite its smaller size, achieving state-of-the-art Inst./ Tool./ ArgN./ and ArgV.(99.51/97.18/96.08/62.10) while substantially improving answer accuracy (45.26). In end-to-end evaluation (Table 4), GPT-4o excels on Op. F1 (66.91) but shows modest Per./Logic. Conversely, OpenEarthAgent delivers balanced gains across Per./Op./Logic./GIS and dominates in tool-order accuracy (AnyOrder/SameOrder/Unique $\approx$ 67-72%), indicating robust trajectory planning rather than isolated tool correctness. Overall, these results suggest that targeted training with tool schemas and trajectories can outperform larger general models on geospatial tool-use, especially under strict order/frequency constraints.
@@ 4.3 Results
Table 4: End-to-end evaluation: (i) F1 scores for tool selection by functional category: Perception (Per.), Operation (Op.), Logic (Logic.), and GIS (GIS); computed against the reference tool sets; (ii) tool-order fidelity: AnyOrder (multiset match, order-agnostic), SameOrder (exact sequence match), and Unique (set match ignoring multiplicity); and (iii) task-level Accuracy, Ans. for non-generative completions and Gen. for image-generation success. GPT-4o shows very high GIS F1, while OpenEarthAgent achieves strong, balanced performance and markedly superior trajectory fidelity, indicating robust multi-step planning rather than isolated tool correctness.
Figure 6: Comparison of model tool-call performance: Open-source models show lower success rates, while GPT-4o and OpenEarthAgent achieve significantly more successful calls.
Figure 7: Category-wise performance: OpenEarthAgent and o4mini lead in index-/ GIS-based scores while some open source perform comparatively on image-based queries.
Tool Call & Success Patterns: Proprietary models (gpt-4o, o4-mini) show strong tool-control, issuing thousands of calls with low error rates ($\sim$7%), indicating stable formatting and reasoning. OpenEarthAgent executes a higher volume of tool calls while achieving a superior overall success rate. In contrast, open-source models such as Qwen2.5-7B, Qwen2.5-3B, and Llama-3.1-8B exhibit high failure rates (43-46%), reflecting weak adherence to the tool schema. Mistral-7B demonstrates a lower error rate but fails to invoke tools at the required adequacy to meet task demands. Despite extensive tool usage, Llama-3.1-8B and Mistral-7B fail to converge to a final answer within the allowed step, resulting in 94% and 98% incomplete task executions, respectively. Overall, performance as shown in Fig. 7 varies by call volume and precision: robust models maintain low error rates, while others struggle with consistent, well-formed tool invocations.
Category Scores: As seen in Fig. 7, OpenEarthAgent (79.43%) and gpt4o (79.39%) achieve the strongest index-based scores, substantially outperforming all baselines. GIS-based tasks reveal a sharper degradation, with OpenEarthAgent (55.77 %) followed by gpt4o (41.95 %) and o4mini (39.55%). Open-source models such as qwen2-5-7b,qwen3-4b, and llama3.1-8b perform comparably well on image-based queries, while smaller models approach zero on index tasks and score only marginally on GIS and image-based tasks. Overall, the results highlight a significant performance gap and show that index- and GIS-based reasoning remains particularly difficult for smaller open models.

### v3

@@ Abstract
Recent progress in multimodal reasoning has enabled agents that interpret imagery, connect it with language, and execute structured analytical tasks. Extending these capabilities to remote sensing remains challenging, as models must reason over spatial scale, geographic structures, and multispectral indices while maintaining coherent multi-step logic. To address this gap, we introduce OpenEarthAgent, a unified framework for tool-augmented geospatial reasoning trained on satellite imagery, natural-language queries, and structured reasoning traces. Beyond serving as a benchmark, OpenEarthAgent establishes a cohesive agentic architecture built around a unified executable tool registry and trajectory-based policy learning. The framework standardizes heterogeneous visual, spectral, GIS, and georeferenced raster operations under a consistent callable schema, enabling modular orchestration and deterministic execution. Training is performed via supervised fine-tuning on structured reasoning trajectories with deterministic replay validation to ensure executability and spatial correctness. The accompanying corpus comprises 14,538 training and 1,169 evaluation instances with over 107K reasoning steps, spanning urban, environmental, disaster, and infrastructure domains and incorporating GIS operations alongside index analyses such as NDVI, NBR, and NDBI. Grounded in explicit reasoning traces, the learned agent demonstrates structured reasoning, stable spatial understanding, and interpretable tool-driven behaviour across diverse EO scenarios. We report consistent improvements over a strong baseline and competitive performance against recent open and closed-source models. Our code and trained models will be publicly available.
@@ 2 Related Work
An emerging wave of EO research is now bringing these agentic principles into geospatial contexts. ThinkGeo [shabbir2025thinkgeo] frames EO QA as tool-augmented reasoning, benchmarking ReAct-style chains and exposing persistent weaknesses in coordinate consistency, spatial grounding, and multi-step planning. Earth-Agent [feng2025earth] broadens the tool ecosystem to include spectral products and standardized interfaces but still relies on largely predefined workflows, limiting physically verifiable GIS and index-based reasoning. Geo-OLM [stamoulis2025geo] explores plan-tool-verify prompting for compact models, improving efficiency but depending primarily on prompt-level heuristics rather than learned adaptive policies. Building on this gap, we introduce an EO-native agentic stack that unifies dataset construction, supervised reasoning alignment, and evaluation under a physically grounded framework. OpenEarthAgent enables broad orchestration of GIS, index-based, optical, and SAR operations through structured instruction-tool pairs; trains on detailed reasoning traces with a dedicated evaluation set; and supports interpretable, verifiable, multi-step reasoning across diverse geospatial conditions.
@@ 3.1 Dataset Curation
Table 2: Comprehensive overview of data sources used for dataset construction, across optical, SAR, GIS, and multispectral domains. These sources provide complementary spatial resolutions and annotations, ranging from urban and transport monitoring to environmental and disaster assessment, forming a rich foundation for multi-sensor, tool-based geospatial reasoning. GSD = ground sample distance, B-Box = bounding box, Seg. = segmentation.
@@
Figure 2: Unified data-curation pipeline for the OpenEarthAgent dataset. RGB/SAR imagery, GIS spatial layers, and index-driven datasets are processed through dedicated filtering (annotation checks, POI constraints, and spectral-change signals such as NDVI/NBR/NDBI). The resulting samples are merged into unified JSON records with imagery metadata and geopackage information. A question synthesizer and conversational generator produce tool-grounded queries and reasoning traces, which are automatically validated to ensure geographic and syntactic correctness before forming the final corpus for agentic training and evaluation.
Figure 3: Representative query-reasoning trajectories from dataset, showing interleaving of natural-language thoughts, tool executions, and observations across tasks such as object localization, direction estimation, spectral mapping, change detection, and GIS-based analysis. Examples are simplified excerpts of longer multi-step reasoning traces.
@@
Dataset Statistics: Fig. 4 summarizes the scale, distribution, and structure of the OpenEarthAgent dataset. The training split contains 14,538 instances, and the benchmark split contains 1,169 evaluation samples, each with multimodal imagery, a natural-language query, and a structured reasoning trace with interleaved tool calls and intermediate observations. The training set contains 100,656 reasoning steps (average 6.92 per query), and the benchmark split contains 7,064 reasoning steps (average 6.04 per query). Reasoning steps correspond to explicit thought-action-observation transitions, while tool calls capture the diversity of operational pathways required for solving curated agentic tasks. The breadth of modalities, thematic coverage, and depth of reasoning traces enable rigorous evaluation of grounded, interpretable, and tool-augmented geospatial reasoning.
@@ 3.2 Methodology
Overview: We present OpenEarthAgent, a unified agentic learning framework for multimodal reasoning over earth observation data. The framework integrates tool-driven interaction and language-grounded reasoning into a single end-to-end formulation. Unlike conventional, single-step multimodal models, OpenEarthAgent decomposes geospatial reasoning into an organized sequence of perception and action steps, where the agent learns to perceive, reason, and act through callable tools operating on visual, textual, and spatial modalities. Given a natural-language instruction $\mathbf{u}$ and corresponding multimodal inputs $\mathbf{v}$ (e.g., RGB imagery, paired change-detection inputs, SAR, GIS layers, or georeferenced raster data), the agent generates multi-step reasoning trajectories that interleave internal thoughts with explicit tool invocations and their returned outputs (see Fig. 5). At each step, the agent maintains a short-term working memory that contains: instructions, observations, prior tool calls, spatial metadata, and contextual feedback from previous executions, which enable iterative reasoning and consistent spatial grounding. Training follows a supervised fine-tuning procedure on validated reasoning traces, improving syntactic validity, spatial consistency, and multi-step reasoning coherence across diverse geospatial contexts.
Unified Tool Registry: To standardize the diverse operators used in EO reasoning, we construct a unified tool registry that abstracts heterogeneous visual, spectral, and GIS operations through a consistent callable schema:
@@
To ensure efficiency and consistency, intermediate outputs (e.g., derived vector layers, raster subsets, index maps, computed geometries) are stored in a tool execution cache memory. This cache includes geospatial archives and vector/raster layers that can be reused by subsequent tool calls within the same trajectory. Caching prevents redundant computation and ensures deterministic replay of reasoning steps. New tools can be integrated by registering their schema $\mathcal{M}_{j}$ without retraining the backbone model, ensuring extensibility while maintaining interface consistency. We outline the tool categories:
a) Perceptual Tools. These operators ground language into image-space entities. TextToBbox, ObjectDetection, and RegionAttributeDescription convert free-form queries into spatial primitives such as bounding boxes and semantic tags. CountGivenObject and SegmentObjectPixels support object counting and pixel-level area estimation. ChangeDetection analyzes multi-date imagery to detect structural changes such as damage or flooding. b) GIS Computation Tools. These tools enable geographic reasoning through geometric and coordinate-aware computation. Operators such as GetAreaBoundary, AddPoisLayer, and ComputeDistance manipulate vector geometries within standard coordinate reference systems. Outputs include numerical measures (e.g., distances, buffers, zonal statistics) or derived layers visualized via DisplayOnMap. c) Spectral Tools. These operators perform band arithmetic and spectral analysis on raster inputs. AddIndexLayer, ComputeIndexChange, and ShowIndexLayer compute and visualize indices such as NDVI, NBR, and NDBI. They enable quantitative reasoning over environmental dynamics. d) Georeferenced Raster Tools. Tools such as GetBboxFromGeotiff and DisplayOnGeotiff operate directly on georeferenced rasters. They support coordinate extraction, bounding box derivation, and visualization aligned with spatial metadata, ensuring spatial consistency. e) Utility Tools. Utilities such as Calculator, Solver, and Plot support arithmetic and symbolic reasoning. Visualization tools DrawBox and AddText display intermediate results. OCR and GoogleSearch extend multimodal understanding through text extraction and contextual retrieval. Terminate signals completion of a reasoning trajectory.
Reasoning Trajectories and Training Objective: Each training sample encodes a structured reasoning trajectory that connects an instruction $\mathbf{u}$, observations $\mathbf{v}$, and a sequence of tool interactions. Formally, the trajectory is represented as:
@@ 4 Experiments
In this section, we outline implementation details, describe the two evaluation modes, define metrics used to quantify reasoning fidelity, tool effectiveness, and geospatial accuracy, and discuss key experimental results and insights.
Implementation Details: OpenEarthAgent is finetuned using Qwen3-4B-Instruct-2507 as the base model. Training was performed on LLM controller on four NVIDIA A100 GPUs (40 GB). The base model is loaded via Unsloth’s FastLanguageModel, enabling efficient multi-GPU distributed training. The model is trained for one epoch with a learning rate of $2\times 10^{-5}$, cosine learning rate scheduling, 0.05 warmup ratio, and a batch size of 16. The maximum sequence length is set to 4096 tokens to support long-context conversational tool interactions. LLM is finetuned on conventional data, where the model alternates between text generation and tool invocation, providing tool outputs and feedback in JSON format, embedded in its user and assistant turns. Training applies response-only masking, computing loss exclusively on assistant-generated tokens, enabling accurate tool calls and structured responses without optimizing on prompt text or external tool outputs. This enables the model to establish a good grounding for tool invocation.
@@ 4.1 Baseline Evaluation
End-to-End Evaluation: This mode tests full autonomous execution with live tool use. The model issues tool calls, forms arguments, and reasons step-by-step based on prior outputs. This captures operational robustness, error handling, and the effectiveness of linking perception to action in geospatial workflows.
@@
Table 4: End-to-end evaluation: (i) F1 scores for tool selection across: Perception (Per.), Operation (Op.), Logic (Logic.), and GIS (GIS.); (ii) tool-order fidelity: AnyOrder (multiset match, order-agnostic), SameOrder (exact sequence match), and Unique (set match, no multiplicity); and (iii) task-level Accuracy, Ans. for non-generative and Gen. for image-generation. GPT-4o attains high GIS F1, while OpenEarthAgent achieves strong, balanced performance with superior trajectory fidelity.
@@
Figure 6: Tool-call performance: Open-source models show lower success rates, while GPT family and OpenEarthAgent achieve substantially higher success rates.
Figure 7: Category-wise performance: OpenEarthAgent and gpt-5 lead in index/ GIS score, while open source performs comparatively on image-based queries.
4.2 Results
In step-by-step evaluation (Table 3), frontier models lead: o4-mini score highest on Summ. (89.48) while gpt4o shows strong Inst. Acc. Among open-source models, Qwen2.5-I (7B) performs competitively (Inst. 94.08, Tool. 85.51), whereas other 7B–8B models degrade, particularly on ArgV. prediction. OpenEarthAgent (4B) substantially narrows this gap despite its smaller size, achieving state-of-the-art Inst./ Tool./ ArgN./ and ArgV.(99.51/97.18/96.08/62.10) while substantially improving Answer Acc. (45.26). In the end-to-end evaluation (Table 4), GPT-5 avoids external logic tools and relies on internal numerical and analytical computations. GPT-4o leads in Op. F1 (66.91) and Gen. Acc. (77.93) but remains modest in Per./Logic, whereas OpenEarthAgent delivers balanced gains across Per./ Op./Logic./GIS and dominates in tool-order accuracy (AnyOr./SameO. /Uni. $\approx$ 67-72%), indicating robust trajectory planning rather than isolated tool correctness. Overall, these results suggest that targeted training with tool schemas and trajectories can outperform larger general purpose models for geospatial tool use, especially under strict order/frequency constraints.
Tool Call & Success Patterns: Proprietary models (gpt-5, gpt-4o and o4-mini) show strong tool-control, issuing thousands of calls with low error rates ($\sim$7%), indicating stable formatting. OpenEarthAgent executes a higher volume of tool calls while achieving a superior overall success rate. In contrast, open-source models such as Qwen2.5-7B, Qwen2.5-3B, and Llama-3.1-8B exhibit high failure rates (43-46%), reflecting weak adherence to the tool schema. Mistral-7B demonstrates a lower error rate but fails to invoke tools at the required adequacy to meet task demands. Despite extensive tool usage, Llama-3.1-8B and Mistral-7B fail to converge to a final answer within the allowed step, resulting in 94% and 98% incomplete task executions, respectively. Overall, performance as shown in Fig. 7 varies by call volume and precision: robust models maintain low error rates, while others struggle with consistent, well-formed tool invocations.
Category Scores: As seen in Fig. 7, OpenEarthAgent (79.43%), gpt-4o (79.39 %), and gpt-5 (78.94%) achieve the strongest index-based scores, substantially outperforming all baselines. GIS-based tasks reveal a sharper degradation, with OpenEarthAgent (55.77 %) followed by gpt-5 (54.94 %), gpt4o (41.95 %), and o4mini (39.55%). Open-source models; qwen2.5-7b, qwen3-4b, and llama3.1-8b perform comparably well on image-based queries, while smaller models approach zero on index tasks and score only marginally on GIS and image-based tasks. Overall, the results highlight a significant performance gap and show that index- and GIS-based reasoning remains difficult for smaller open models.
Additional analyses and ablations are provided in the supplementary material.